# Customized SAM 2 for Referring Remote Sensing Image Segmentation

### v1

@@ 2 Related work
Segment Anything Model. SAM [10] is an interactive model designed to generate non-semantic segmentation masks based on a variety of input prompts. Leveraging large-scale training data, it demonstrates strong generalization across diverse segmentation tasks. SAM has also been widely utilized in multiple domains, including remote sensing image analysis [28], video object tracking [2, 34], and medical image segmentation [39, 3]. To improve its accuracy and efficiency, several optimized versions [41, 32, 43] have been developed. More recently, SAM 2 [24] has been introduced for video segmentation [25, 12], incorporating multi-scale feature encoding to enhance segmentation robustness. Despite these advances, SAM remains limited by its lack of linguistic understanding, preventing it from directly handling referring segmentation tasks that require textual guidance. To address this limitation, recent works [11, 33, 23] have explored the integration of multimodal large language models (MLLMs) to enhance SAM’s ability to process language-based instructions. LISA [11], for example, fine-tunes LLaVA [16] to derive hidden embeddings, thereby producing multimodal features for enhanced segmentation. u-LLaVA [33] extends this paradigm by enabling simultaneous region- and pixel-level segmentation through multi-task learning. Due to its lightweight pre-fusion strategy, EVF-SAM [42] leverages joint visual-language encoding to enhance text-driven segmentation prompts, thereby achieving superior segmentation accuracy. However, while these models have demonstrated strong performance on natural images, their effectiveness in complex remote sensing scenes remains limited, highlighting the need for further refinement and adaptation.
@@ 3.3 Bidirectional Hierarchical Fusion Module
As illustrated in Fig. 3, our bidirectional hierarchical fusion module begins with the preprocessed SAM 2 image feature $F_{i}$ of the current layer being projected to a lower dimensionality through a linear layer and an activation function. Simultaneously, the text feature $T_{i}$ of the current layer is also projected to match the dimensionality of the image feature using a linear layer. Subsequently, the image and text features are utilized as query embeddings and key-value pairs, respectively, and undergo cross-attention interaction to capture modality-specific dependencies. The resulting features are then integrated with the pre-interaction representations through element-wise addition. Following this, linear layers are used to restore the dimensionality of both the image and text features. The above process can be expressed by the following equation:
$F_{i}^{{}^{\prime}}=\sigma(\text{Linear}(F_{i})),T_{i}^{{}^{\prime}}=\text{Linear}(T_{i}),$ (2)
$F_{i}^{{}^{\prime\prime}}=\text{MHCA}(F_{i}^{{}^{\prime}},T_{i}^{{}^{\prime}})+F^{{}^{\prime}}_{i},$
$T_{i}^{{}^{\prime\prime}}=\text{MHCA}(T_{i}^{{}^{\prime}},F_{i}^{{}^{\prime}})+T^{{}^{\prime}}_{i},$
where $\sigma$ denotes the activation function GeLU and MHCA represents the multi-head cross-attention layer.
To preserve textual integrity during visual enhancement, the text feature is weighted and summed with its pre-interaction representation. Meanwhile, after skip-connection, the visual feature is summed with the original image feature $F_{in}$ of the current layer, then separately processed by the MLP branch and the linear branch, followed by weighted fusion. The entire process is depicted as follows:
@@
$F_{out}=$ $F^{{}^{\prime\prime\prime}}_{i}+\text{MLP}(\text{LN}(F^{{}^{\prime\prime\prime}}_{i}))+$
$\alpha_{i}\text{Linear}(\sigma(\text{Linear}(F^{{}^{\prime\prime\prime}}_{i}))),$
@@
After feature encoding, the original text feature $T$ encoded by the union encoder is used to further guide the visual feature $F$ at the high-level. Specifically, the visual feature acts as query and interacts with the text feature as key-value pairs through multi-head cross-attention. The resulting feature is then element-wise multiplied with the visual feature before interaction. The process is as shown in the following equation:
$F_{en}=\text{MHCA}(F,T)*F,$ (4)
subsequently, the obtained text-guided hierarchical feature $F_{en}$ is input into the SAM 2 decoder for accurate decoding.
@@ 3.4 Mask Prompt Generator
As illustrated in Fig. 2, we first employ the multimodal [CLS] token $V_{cls}$ as query and the visual embeddings $V$ as key-value pairs for cross-attention computation. The interaction result is then element-wise multiplied with the multimodal [CLS] token to further align the multimodal token with visual information. Subsequently, the visual embeddings are reshaped into feature maps with dimensions $\frac{H_{u}}{p}\times\frac{W_{u}}{p}$, and the multimodal token is passed through a linear layer and broadcasted to match the shape of the visual embeddings. After element-wise multiplication, the result is input into the mask generator, composed of MLP layers, to produce the pseudo-mask $M_{p}\in\mathbb{R}^{\frac{H_{u}}{p}\times\frac{W_{u}}{p}}$ that is well-aligned with multimodal information. The formal description of this process is provided as follows:
$V_{cls}^{{}^{\prime}}=\text{MHCA}(V_{cls},V),$ (5)
$V^{{}^{\prime}}=R(V)*\text{Linear}(V_{cls}^{{}^{\prime}}),$
$M_{p}=\text{MLP}(V^{{}^{\prime}}),$
where MHCA represents the multi-head cross-attention layer, and $R(\cdot)$ denotes the reshape operation.
@@ 3.6 Training Loss
$\nabla_{pre}=Absd_{h}(M_{pre})+Absd_{v}(M_{pre})$ (7)
$\nabla_{gt}=Absd_{h}(M_{gt})+Absd_{v}(M_{gt})$
$\mathcal{L}_{tbl}=\frac{1}{N}\sum_{i=1}^{N}(\text{Linear}(T_{s})({\nabla_{pre}}_{i}-{\nabla_{gt}}_{i}))^{2}$
@@ 4.3 Comparison with State-of-the-Art Methods
RefSegRS dataset. We conduct a comprehensive comparison between our method, RS2-SAM 2, and several state-of-the-art approaches on the RefSegRS [38] dataset. The results of the comparison are reported in Tab. 1. It can be observed that our RS2-SAM 2 achieves 88.03% oIoU and 85.21% mIoU on the validation set, surpassing the previous best method, RMSIN, by 5.62% in oIoU and 11.37% in mIoU. On the test set, it reaches 80.87% oIoU and 73.90% mIoU, outperforming FIANet [14] by 2.55% and 5.23%, respectively. Moreover, the Pr metric exhibits significant improvements across all thresholds, particularly at Pr@0.7, Pr@0.8, and Pr@0.9, demonstrating the model’s strong multimodal segmentation capabilities.
RRSIS-D dataset. We also conduct comparative experiments between our RS2-SAM 2 and existing methods such as LGCE [38], LAVT [35], EVF-SAM [42] and RMSIN [19], etc., on the RRSIS-D [19] dataset, with results documented in Tab. 2. On this dataset, our method achieves the best performance, surpassing the current state-of-the-art method RMSIN on the validation set by 4.59% in Pr@0.5, 5.86% in Pr@0.6, 6.44% in Pr@0.7, 5.28% in Pr@0.8, 5.97% in Pr@0.9, 1.89% in oIoU, and 3.71% in mIoU. It also significantly outperformed existing methods on the test set, demonstrating its strong capability in remote sensing object segmentation.
@@ 4.4 Model Analysis
Components Analysis. To investigate the impact of key components in our model, we first construct a baseline model consisting solely of SAM 2 and the union encoder. As shown in Tab. 3, when the text-guided boundary loss ($\mathcal{L}_{tbl}$) is incorporated during training, the model achieves 38.63% mIoU and 57.36% oIoU, surpassing the baseline by 1.99% and 1.85%, respectively. Subsequently, on top of the model with $\mathcal{L}_{tbl}$, we add the mask prompt generator (MPG), which results in the RS2-SAM 2 achieving 60.20% mIoU and 70.89% oIoU, an improvement of 21.57% and 13.53% over the previous model. When the bidirectional hierarchical fusion module (BHFM) is added to the baseline model with $\mathcal{L}_{tbl}$, the model shows a further increase of 30.08% in mIoU and 21.00% in oIoU, further validating the superiority of this module. Finally, when all components are integrated, our RS2-SAM 2 achieves the best performance, with a mIoU of 73.90% and an oIoU of 80.87%.
@@
Mask Prompt Generator. In this section, we explore different interaction forms within the mask prompt generator (MPG), with results presented in Tab. 4. When the multi-head cross-attention (MHCA) between the multimodal token $V_{cls}$ and visual embeddings $V$ is omitted, the performance of RS2-SAM 2 drops by 2.31% in mIoU and 0.98% in oIoU. These findings underscore the importance of further strengthening the semantic connections between the multimodal token and visual embeddings during the multimodal mask prompt generation process.
Bidirectional Hierarchical Fusion Module. The impact of different settings of the bidirectional hierarchical fusion module (BHFM) is also noteworthy. First, we examine how the structure of the BHFM layer during feature encoding affects the model’s performance. One simple approach is to use a linear layer structure similar to the adapter, without interacting with text feature, referred to as “Linear”. Another approach is to enhance visual feature using only text feature, without back-propagating visual information, referred to as “Uni”. The third approach is bidirectional enhancement of both text and visual features, referred to as “Bi”. The experimental results in Tab. 4 show that the layer-by-layer bidirectional enhancement of text and visual features is more beneficial for imparting linguistic information to the visual feature. This enables refined linguistic enhancement of the visual feature from low-level to high-level, leading to more accurate segmentation of remote sensing objects.
Additionally, we investigate the impact of BHFM components on segmentation performance by setting up experiments where the BHFM cross-attention (BC) after encoding and the BHFM layer (BL) without the encoder are removed. The results in Tab. 4 indicate that both interactions during and after encoding are essential. The hierarchical interaction combining both can help the model understand text features from a global to a local perspective, enhancing its pixel-level understanding of text.

### v2

@@ 2 Related work
Segment Anything Model. SAM [10] is an interactive model designed to generate non-semantic segmentation masks based on a variety of input prompts. Leveraging large-scale training data, it demonstrates strong generalization across diverse segmentation tasks. SAM has also been widely utilized in multiple domains, including remote sensing image analysis [28], video object tracking [2, 34], and medical image segmentation [39, 3]. To improve its accuracy and efficiency, several optimized versions [41, 32, 44] have been developed. More recently, SAM 2 [24] has been introduced for video segmentation [25, 12], incorporating multi-scale feature encoding to enhance segmentation robustness. Despite these advances, SAM remains limited by its lack of linguistic understanding, preventing it from directly handling referring segmentation tasks that require textual guidance. To address this limitation, recent works [11, 33, 23, 42] have explored the integration of multimodal large language models (MLLMs) to enhance SAM’s ability to process language-based instructions. LISA [11], for example, fine-tunes LLaVA [16] to derive hidden embeddings, thereby producing multimodal features for enhanced segmentation. u-LLaVA [33] extends this paradigm by enabling simultaneous region- and pixel-level segmentation through multi-task learning. Due to its lightweight pre-fusion strategy, EVF-SAM [43] leverages joint visual-language encoding to enhance text-driven segmentation prompts, thereby achieving superior segmentation accuracy. However, while these models have demonstrated strong performance on natural images, their effectiveness in complex remote sensing scenes remains limited, highlighting the need for further refinement and adaptation.
@@ 3.3 Bidirectional Hierarchical Fusion Module
As illustrated in fig. 3, our bidirectional hierarchical fusion module begins with the preprocessed SAM 2 image feature $F_{i}$ of the current layer being projected to a lower dimensionality through a linear layer and an activation function, yielding $F_{i}^{{}^{\prime}}$. Simultaneously, the text feature of the current layer, $T_{i}$, is also projected via a linear layer to obtain $T_{i}^{{}^{\prime}}$, ensuring dimensional consistency with the image feature. Subsequently, the image and text features are utilized as query embeddings and key-value pairs, respectively, and undergo cross-attention interaction to capture modality-specific dependencies. The resulting features are then integrated with the pre-interaction representations through element-wise addition. Following this, linear layers are used to restore the dimensionality of both the image and text features, producing the updated representations $F_{i}^{{}^{\prime\prime}}$ and $T_{i}^{{}^{\prime\prime}}$.
To preserve textual integrity during visual enhancement, the text feature is weighted and summed with its pre-interaction representation. Meanwhile, after skip-connection, the visual feature is summed with the original image feature $F_{in}$ of the current layer, then separately processed by the MLP branch and the linear branch, followed by weighted fusion. The above process can be expressed by the following equation:
@@
$F_{out}=$ $F^{{}^{\prime\prime\prime}}_{i}+\text{MLP}(\text{LN}(F^{{}^{\prime\prime\prime}}_{i}))+\alpha_{i}\text{Linear}(\sigma(\text{Linear}(F^{{}^{\prime\prime\prime}}_{i}))),$
@@
After feature encoding, the original text feature $T$ encoded by the union encoder is used to further guide the visual feature $F$ at the high-level. Specifically, the visual feature acts as query and interacts with the text feature as key-value pairs through multi-head cross-attention. The resulting feature is then element-wise multiplied with the visual feature before interaction, yielding the text-guided hierarchical feature $F_{en}$, which is subsequently fed into the SAM 2 decoder for accurate decoding.
@@ 3.4 Mask Prompt Generator
As illustrated in fig. 2, we first employ the multimodal [CLS] token $V_{cls}$ as query and the visual embeddings $V$ as key-value pairs for cross-attention computation. The interaction result is then element-wise multiplied with the multimodal [CLS] token to further align the multimodal token with visual information. Subsequently, the visual embeddings are reshaped into feature maps with dimensions $\frac{H_{u}}{p}\times\frac{W_{u}}{p}$, and the multimodal token is passed through a linear layer and broadcasted to match the shape of the visual embeddings. After element-wise multiplication, the result is input into the mask generator, composed of MLP layers, to produce the pseudo-mask $M_{p}\in\mathbb{R}^{\frac{H_{u}}{p}\times\frac{W_{u}}{p}}$ that is well-aligned with multimodal information.
@@ 3.6 Training Loss
$\nabla_{pre}=Absd_{h}(M_{pre})+Absd_{v}(M_{pre}),$ (4)
$\nabla_{gt}=Absd_{h}(M_{gt})+Absd_{v}(M_{gt}),$
$\mathcal{L}_{tbl}=\frac{1}{N}\sum_{i=1}^{N}(\text{Linear}(T_{s})({\nabla_{pre}}_{i}-{\nabla_{gt}}_{i}))^{2},$
@@ 4.3 Comparison with State-of-the-Art Methods
RefSegRS dataset. We conduct a comprehensive comparison between our method, RS2-SAM 2, and several state-of-the-art approaches on the RefSegRS [38] dataset. The results of the comparison are reported in table 1. It can be observed that our RS2-SAM 2 achieves 88.03% oIoU and 85.21% mIoU on the validation set, surpassing the previous best method, RMSIN, by 5.62% in oIoU and 11.37% in mIoU. On the test set, it reaches 80.87% oIoU and 73.90% mIoU, outperforming FIANet [14] by 2.55% and 5.23%, respectively. Moreover, the Pr metric exhibits significant improvements across all thresholds, particularly at Pr@0.7, Pr@0.8, and Pr@0.9, demonstrating the model’s strong multimodal segmentation capabilities.
RRSIS-D dataset. We also conduct comparative experiments between our RS2-SAM 2 and existing methods such as LGCE [38], LAVT [35], EVF-SAM [43] and RMSIN [19], etc., on the RRSIS-D [19] dataset, with results documented in table 2. On this dataset, our method achieves the best performance, surpassing the current state-of-the-art method RMSIN on the validation set by 4.59% in Pr@0.5, 5.86% in Pr@0.6, 6.44% in Pr@0.7, 5.28% in Pr@0.8, 5.97% in Pr@0.9, 1.89% in oIoU, and 3.71% in mIoU. It also significantly outperformed existing methods on the test set, demonstrating its strong capability in remote sensing object segmentation.
@@
4.4 Model Analysis
Components Analysis. To investigate the impact of key components in our model, we first construct a baseline model consisting solely of SAM 2 and the union encoder. As shown in table 3, when the text-guided boundary loss ($\mathcal{L}_{tbl}$) is incorporated during training, the model achieves 38.63% mIoU and 57.36% oIoU, surpassing the baseline by 1.99% and 1.85%, respectively. Subsequently, on top of the model with $\mathcal{L}_{tbl}$, we add the mask prompt generator (MPG), which results in the RS2-SAM 2 achieving 60.20% mIoU and 70.89% oIoU, an improvement of 21.57% and 13.53% over the previous model. When the bidirectional hierarchical fusion module (BHFM) is added to the baseline model with $\mathcal{L}_{tbl}$, the model shows a further increase of 30.08% in mIoU and 21.00% in oIoU, further validating the superiority of this module. Finally, when all components are integrated, our RS2-SAM 2 achieves the best performance, with a mIoU of 73.90% and an oIoU of 80.87%.
Mask Prompt Generator. In this section, we explore different interaction forms within the mask prompt generator (MPG), with results presented in table 4. When the multi-head cross-attention (MHCA) between the multimodal token $V_{cls}$ and visual embeddings $V$ is omitted, the performance of RS2-SAM 2 drops by 2.31% in mIoU and 0.98% in oIoU. These findings underscore the importance of further strengthening the semantic connections between the multimodal token and visual embeddings during the multimodal mask prompt generation process.
@@ 4.4 Model Analysis
Bidirectional Hierarchical Fusion Module. The impact of different settings of the bidirectional hierarchical fusion module (BHFM) is also noteworthy. First, we examine how the structure of the BHFM layer during feature encoding affects the model’s performance. One simple approach is to use a linear layer structure similar to the adapter, without interacting with text feature, referred to as “Linear". Another approach is to enhance visual feature using only text feature, without back-propagating visual information, referred to as “Uni". The third approach is bidirectional enhancement of both text and visual features, referred to as “Bi". The experimental results in table 4 show that the layer-by-layer bidirectional enhancement of text and visual features is more beneficial for imparting linguistic information to the visual feature. This enables refined linguistic enhancement of the visual feature from low-level to high-level, leading to more accurate segmentation of remote sensing objects.
Additionally, we investigate the impact of BHFM components on segmentation performance by setting up experiments where the BHFM cross-attention (BC) after encoding and the BHFM layer (BL) without the encoder are removed. Results in table 4 indicate that both interactions during and after encoding are essential. The hierarchical interaction helps the model understand text features from a global to local perspective, enhancing its pixel-level understanding of text.